At some time in the future, John will be 38 years old. At that time, he will be three times as old as Sally. Sally is now 7 years old. How old is John now
Answer by macston(5194)

When John is 38, he will be 3 times as old as Sally, so Sally will be 38/3=12 2/3 years old when John is 38 years old.
Sally is now 7, so that will be (12 2/3-7=5 2/3) 5 and 2/3 years from now.
Since John will then be 38, he is presently (38-5 2/3=32 1/3) 32 1/3 years old.
ANSWER: John is currently 32 1/3 years old